OpenStudy (anonymous):

how do we graph this ?? Any help would be awesome! 16 less than or equal to x2+ y2 less than or equal to 25

myininaya (myininaya):

x^2+y^2=r^2 is a circle remember?

OpenStudy (anonymous):

a donut centered at the origin inner radius 4 outer one 5

OpenStudy (anonymous):

maybe torus?

OpenStudy (anonymous):

spare tire?

myininaya (myininaya):

i was fixing to say that and then shade on the donut

OpenStudy (anonymous):

concentric circles?

myininaya (myininaya):

do not shade the donut hole

OpenStudy (anonymous):

no save that for me!

myininaya (myininaya):

just the donut

OpenStudy (anonymous):

awesome how do you know its a circle

myininaya (myininaya):

because (x-h)^2+(y-k)^2=r^2 is the general form for the a circle

OpenStudy (anonymous):

pythagoras. \[x^2+y^2=r^2\] says distance between (0,0) and (x,y) is r